# Barrow holographic dark energy with varying exponent

### v1

@@ 4.5 Trigonometric functions case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}(2-\alpha-\beta\sin{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\cos{z}}{\alpha+\beta\sin{z}-2}}+\alpha+\beta\sin{z}+1,\hskip 32.00002pt$ (4.10)
$\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\sin{z}+1}{3}+\frac{(\alpha+\beta\sin{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}$
$\ \ \ \ \ \ +\frac{(1+z)\beta\cos{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\sin{z}}}}.$ (4.11)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}(2-\alpha-\beta\cos{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\sin{z}}{\alpha+\beta\cos{z}-2}}+\alpha+\beta\cos{z}+1,\hskip 28.00006pt$ (4.12)
$\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\cos{z}+1}{3}+\frac{(\alpha+\beta\cos{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}$
$\ \ \ \ \ \ \,-\frac{(1+z)\beta\sin{z}}{3}\log{\left[\frac{Q(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\cos{z}}}}$ (4.13)
@@ 5 Conclusions
Nevertheless, the most suitable ansatz for the purpose of Barrow holographic dark energy with varying exponent is the hyperbolic tangent one, since in this case we can easily bound Barrow exponent between 0 and 1 for all redshifts, obtaining a scenario where $\Delta(z)$ is 1 at early times while it becomes 0 at intermediate and late times, as well as in the future. This scenario can also recover the thermal history of the Universe, and the recent transition to acceleration. Furthermore, we investigated the effect of the single parameter $\gamma$ on the dark-energy equation-of-state parameter. We showed that for increasing $\gamma$ the value of $w_{DE}$ decreases, while for some cases it can experience the phantom-divide crossing, and that in the future $w_{DE}$ can either tend to $-1$ or start increasing again. Lastly, for completeness, we also presented the case where the running Barrow exponent has a trigonometric form.
All these features reveal that Barrow holographic dark energy with varying exponent is not only theoretically more justified than the standard, constant-exponent case, but it leads to richer cosmological behavior too. It would be interesting to perform observational confrontation with data from Supernovae type Ia (SNIa), Baryonic Acoustic Oscillations (BAO), Cosmic Microwave Background (CMB), and Cosmic Chronometers (CC) probes, in order to extract constraints on the involved parameters, and apply information criteria in order to compare the scenario with the concordance $\Lambda$CDM cosmology. Such a full observational analysis is left for a future project.

### v2

@@ 4.5 Trigonometric functions case
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}(2-\alpha-\beta\sin{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \,\ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{-(1+z)\beta\cos{z}}{\alpha+\beta\sin{z}-2}}+\alpha+\beta\sin{z}+1,\hskip 32.00002pt$ (4.12)
$\!\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\sin{z}+1}{3}+\frac{(\alpha+\beta\sin{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\sin{z}}{2(\alpha+\beta\sin{z}-2)}}$
$\ \ \ \ \ \ +\frac{(1+z)\beta\cos{z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\sin{z}}}}.$ (4.13)
@@
$\!\!\!\!\!\!\!\!\!\!\!\!\!\frac{-(1+z)\Omega_{DE}^{\prime}}{\Omega_{DE}\left(1-\Omega_{DE}\right)}=\sqrt{\Omega_{DE}}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}(2-\alpha-\beta\cos{z})$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ +\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]}^{\frac{(1+z)\beta\sin{z}}{\alpha+\beta\cos{z}-2}}+\alpha+\beta\cos{z}+1,\hskip 28.00006pt$ (4.14)
$\!\!\!\!\!\!\!w_{DE}=-\frac{\alpha+\beta\cos{z}+1}{3}+\frac{(\alpha+\beta\cos{z}-2)\sqrt{\Omega_{DE}}}{3}\left(\frac{{C}}{3M_{p}^{2}}\right)^{-\frac{1}{2}}\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{\alpha+\beta\cos{z}}{2(\alpha+\beta\cos{z}-2)}}$
$\ \ \ \ \ \ \,-\frac{(1+z)\beta\sin{z}}{3}\log{\left[\frac{P(1-\Omega_{DE})}{\Omega_{DE}}\right]^{\frac{1}{2-\alpha-\beta\cos{z}}}}$ (4.15)
@@ 5 Conclusions
Nevertheless, the most suitable ansatz for the purpose of Barrow holographic dark energy with varying exponent is the hyperbolic tangent one, since in this case we can easily bound Barrow exponent between 0 and 1 for all redshifts, obtaining a scenario where $\Delta(z)$ is 1 at early times while it becomes 0 at intermediate and late times, as well as in the future. This scenario can also recover the thermal history of the Universe, and the recent transition to acceleration. Furthermore, we investigated the effect of the single parameter $\gamma$ on the dark-energy equation-of-state parameter. We showed that for increasing $\gamma$ the value of $w_{DE}$ decreases, while for some cases it can experience the phantom-divide crossing, and that in the future $w_{DE}$ can either tend to $-1$ or start increasing again. For completeness, we performed a basic confrontation of the hyperbolic tangent model against the Supernovae type Ia (SnIa) and Cosmic Chronometers (CC) datasets, as a first evidence that they are viable and consistent with observations. Lastly, we also presented the case where the running Barrow exponent has a trigonometric form.
All these features reveal that Barrow holographic dark energy with varying exponent is not only theoretically more justified than the standard, constant-exponent case, but it leads to richer cosmological behavior too. It would be interesting to perform a full observational confrontation with data from Supernovae type Ia (SNIa), Baryonic Acoustic Oscillations (BAO), Cosmic Microwave Background (CMB), and Cosmic Chronometers (CC) probes, in order to construct the likelihood contours and extract constraints on the involved parameters, and apply information criteria in order to compare the scenario with the concordance $\Lambda$CDM cosmology. Such a full observational analysis is left for a future project.